OpenStudy (anonymous):

simplify (8.37 x 10^16)÷(9 x 10^9) and write the answer in scientific notation.

OpenStudy (anonymous):

837*10^14/9*10^9 =93*10^5 =9.3*10^6

OpenStudy (anonymous):

is it 93 x 10^24

OpenStudy (anonymous):

cause i put 9.3 on the test and i got it wrong so it must be 93

OpenStudy (anonymous):

or i got half credit but not fully wrong

OpenStudy (anonymous):

It's 9.3*10^8

OpenStudy (anonymous):

In scientific notation u dnt write 93 bt 9.3